# Creative Writers’ Attitudes on Writing as Training Data for Large Language Models

### v1

@@ 6. Discussion
As seen in the above quote, writers often traversed a range of emotions in response to the implications of LLMs, and many writers noted that they were not sure what should be done. Here, we turn first to research that has tried to envision more responsible data collection practices. We consider how our findings do and don’t align with archival practices, and how findings from research on archival practices could shed light on how collecting creative writing as training data may be done responsibility. Then, we consider how LLMs disrupt our notion of the creative chain, discuss how LLMs could look more like libraries (which have a well understood and well respected place in the creative chain), and ways in which they may fail to do so. Finally, we propose future research to support these visions.
6.1. Archival Practices for Responsible Data Collection
Recent work compares data collection for machine learning use cases to archival practices (Jo and Gebru, 2020; Desai et al., 2024). Here, we draw on this work and related archival scholarship to briefly discuss relevant concepts and consider their application to creative writing as training data for LLMs.
6.1.1. Community archives
Drawing on scholarship on community archives (e.g. (Flinn et al., 2009; Flinn, 2007; Huvila, 2008; Caswell et al., 2016)), researchers advocate for community contributed training datasets for machine learning, which give groups agency to represent themselves and collectively understand and mitigate the risks of contributing their data (Jo and Gebru, 2020; Desai et al., 2024). We found in our interviews that, for the most part, writers were most interested in participating in data collection projects led by and for other writers.
6.1.2. Access and collection policies
Jo and Gebru (2020) recommend allowing participants to specify access limits when contributing data to community datasets. In line with this recommendation, our interviewees also wanted to have more finely grained control over what their data would be used for. Creative writing community datasets for LLMs should, therefore, have clear mission statements and policies around how data can be used. In addition, many interviewees wanted access to their data to be renegotiated periodically. Given the rapidly changing nature of LLMs, we would recommend time-boxed projects that allow participants to renegotiate their participation every, say, three years.
6.1.3. Transparency
A variety of work has drawn on principles from archives and information science to advocate for better documentation of machine learning datasets (Gebru et al., 2021; McMillan-Major et al., 2021; Bender and Friedman, 2018; Hutchinson et al., 2021). In addition to documenting the contents of a dataset, Desai et al. (2024) note that archives also record the process of data collection. In our own interviews, we found that writers wanted more transparency into training data in order to credit those who contributed. Transparency into the collection process likely would also create good will in participants, who would want to know exactly what they’re signing up for. Furthermore, we advocate for transparency into the use of the resulting models, not just the data collection itself. We imagine something like a yearly shareholder report, where details about model usage are collected and shared with those who contributed to the dataset.
6.1.4. Longevity
As Flinn et al. (2009) show, many community archives are driven by a small set of founding members, and long-term sustainability is challenging after the original driving forces move away or pass on. In the context of training data, we would argue that this challenge is increased by the changing nature of what training data is used for, and difficulties in predicting future use cases. Flinn et al. note that there is sometimes a conflict between autonomy, communities managing their own archives, and sustainability, where outside institutions can help preserve archives over the long term. As proposed above, time-boxed projects that require participants to update their consent could allow projects to transition to new custodians while ensuring participants remain aware of and consenting to such changes.
6.2. Navigating the Role of Training Data in the Creative Chain
As we reported in the creative chain (5.1.1), writers expect their writing to be part of a larger cultural project over which they have little control, and take pleasure in inspiring other writers or contributing to education and literacy projects. However, LLMs challenged their notion of what it meant to participate in the creative chain. While typical engagement with their writing comes directly from other people, for instance another writer mimicking their style, writing fanfiction about their work, or adapting their work to another medium, LLMs create an intermediary. It is unclear how the user of an LLM engages with the writing it was trained on.
@@ 6.2. Navigating the Role of Training Data in the Creative Chain
Some writers envisioned how LLMs could be more aligned with the kinds of cultural projects that they wanted to contribute to. For instance, several writers discussed the positive values of libraries, and considered a world where LLMs were more like libraries. Yet, libraries and language models have important differences. The understood intent of libraries is that a wider variety of people can have direct access to your writing, whereas the contribution of individual authors to an LLM is obscured to the final audience of an LLM output. Libraries are non-profit institutions that serve the general public. They are also physical locations that serve their local community in a variety of ways. Some of these differences could be addressed: LLMs could be created by non-profit institutions to serve the general public. The contributions of individual authors could be made more clear, for instance through work on attributing generated text to specific training data (Deng et al., 2024; Hammoudeh and Lowd, 2024). However, it is likely difficult to completely remove the obscurity that lies between training data and generated text, and this disjuncture may continue to cause unique problems.
@@
6.3. Future Work
We outline two avenues of future work. The first is how to prevent the use of writing as training data without consent. The second is how to create language models that writers would want to contribute to. We see these as the negative and positive sides of futuring: how discourage the future we don’t want, and how do we create a future we desire.
6.3.1. Preventing nonconsensual usage.
While there is preliminary work on detecting if certain text has been used to train a model, or applying watermarks to text, more robust and easy-to-access methods would allow writers to determine if their writing has been used without their consent, which is the first step in recourse. Legal protections could enforce this, but social norms could also discourage institutions from nonconsensual usage. Research into the economic impacts of LLMs could also support this agenda.
6.3.2. Creating writer-friendly language models.
Community datasets would allow writers to control the writing they would like to see contributed only to projects they believed in. Technical work that supports smaller models that can be run locally, or models that can more easily interact with sensitive training data (like (Min et al., 2023)), could support community models which may not have the compute power of corporate ones.
@@ Appendix A Self-Reported Demographic Data from Interviewed Writers
Table 1. Primary Genres of Writers Interviewed
@@
Accordingly, the writers we spoke with also share their work across a variety of publishing or delivery methods, detailed in Table 2. As with genre, the majority of writers we spoke with (18 out of 33) also share their writing through multiple venues and means.
Table 2. Publishing or Distribution Models for Writers Interviewed
@@
Table 3. Compensation for Creative Work for Writers Interviewed
@@
Majority of income comes from creative writing practice 11
Majority of income comes from other writing-intensive profession 10
Majority of income does not come from writing 8
@@
We also noted how writers self-reported on their use of LLMs, see Table 4. The vast majority (23 out of 33) were ”dabblers”—those who had tried out LLMs, but do not use them in a regular or intensive manner. A significant minority (8 out of 33) use LLMs regularly in either their paratextual or central creative writing practices.
@@ B.3. Knowledge of and attitudes towards language model training data
    • Note: Go through these questions for each in stimulus 2

### v2

@@ 6. Discussion
As seen in the above quote, writers often traversed a range of emotions in response to the implications of LLMs, and many writers noted that they were not sure what should be done. Overall, while writers had concerns about the implications of language models abstractly, most writers expressed distaste for large corporations using their work for dubious or unclear ends, and making a profit along the way.
An implicit theme in many of our interviews was the way in which LLMs turned writing into data. Writers consider their work—writing—to be valued in a variety of ways (e.g. appreciation, compensation, prestige) by a human readership. LLMs, in defiance of these norms, subsume writing into the larger and more contentious category of data, where value becomes murky and the audience becomes algorithmic rather than human. Data has been analogized to labor (Arrieta-Ibarra et al., 2018), property (Prainsack, 2019), and natural resource (e.g. ”data mining”). Such analogies, although never perfect, attempt to sort through how data should be valued, especially data generated by, and often about, people. But is writing data?
We argue that writing becomes data when it is used to train generative models, and this transformation creates the issues writers raise, many of which stem from a reorientation of their audience from people to algorithms or perhaps companies. Writers noted the power differential between themselves and the corporations making use of their work. This can be understood through the frame of precarious work (Allan et al., 2021); as writers feel that their work is increasingly precarious, the power differential between themselves and the organizations seeking to train LLMs grows larger. [^1] Such power differentials have been explored in fan fiction communities. For instance, Busse and Farley (2013) report on fan fiction writers who note that the balance of power between two fan writers and between a fan writer and a professional writer is completely different, resulting in different social norms for remixing professional work than fan work.
We can also see this dynamic as a form of extraction, in which writing is ”taken” from writers for the profit of others. Couldry and Mejias (2019) propose ‘data colonialism’ as a new form of colonialism to make sense of the use of large amounts of data by a small group of corporate and government actors. If data colonialism is a useful framework to understand data collection for language models, which we think it may be, then we can turn to theories of decolonialism to understand how to resist it. Mohamed et al. (2020) put forth the idea of dismantling power assymmetries to resist data colonialism. This moves beyond our initial exploration of how data may or may not be collected in this context.
While our findings often align with those studying visual artist communities—work by Jiang et al. (2023) finds that AI image generators create a chilling effect on cultural production, similar to our findings about industry impacts on writers—the direct economic impacts seem to be distinct (e.g. few of our interviewees expected to see personal economic impacts). However, the experience of precarious work and a lack of power remain the same. We believe that decolonial thinking can address a multitude of concerns (e.g. through pro-worker AI policies (Acemoglu et al., 2023) or public AI initiatives (Marda et al., 2024)) even as they differ in their specifics across communities.
When asked who they expect to best advocate for their interests, writers did not think governments, publishers, universities, and certainly not technology companies, had their interests at heart. Some writers relied on their agents, but most writers would only trust other writers to safely and respectfully steer LLM creation. We recommend that writers form political communities which can take the helm in determining how their work is used.
@@
As we reported in the creative chain (5.1.1), while writers expect their writing to be part of a larger cultural project over which they have little control, LLMs challenged their notion of what it meant to participate in the creative chain. Some writers envisioned how LLMs could be more aligned with the kinds of cultural projects that they wanted to contribute to. For instance, several writers discussed the positive values of libraries, and considered a world where LLMs were more like libraries. Yet, libraries and language models have important differences. The understood intent of libraries is that a wider variety of people can have direct access to your writing, whereas the final audience of an LLM may be ignorant of individual author contributions. Libraries are non-profit institutions that serve the general public. They are also physical locations that serve their local community in a variety of ways.
Some of these differences could be addressed: LLMs could be created by non-profit institutions to serve the general public, as recommended in (Marda et al., 2024) and seen in projects like (Le Scao et al., 2023) and (Soldaini et al., 2024). The contributions of individual authors could be made more clear, for instance through work on attributing generated text to specific training data (Deng et al., 2024; Hammoudeh and Lowd, 2024). However, it is likely difficult to completely remove the obscurity that lies between training data and generated text, and this disjuncture may continue to cause unique problems.
@@
7. Future Work
We outline two avenues of future work. The first is how to prevent the use of writing as training data without consent. The second is how to create models that writers would want to contribute to. We see these as the negative and positive sides of futuring: how to discourage a future we don’t want, and create a future we desire.
Preventing nonconsensual usage. While there is preliminary work on detecting if certain text has been used to train a model (Hammoudeh and Lowd, 2024), or applying watermarks to text (Lau et al., 2024; Qiang et al., 2023; Yoo et al., 2023), more robust and easy-to-access methods would allow writers to determine whether their writing has been used without their consent, which is the first step in recourse. Legal protections could enforce this, but social norms could also discourage institutions from nonconsensual usage. [^2] For example, fan fiction communities have been found to have highly consistent social norms which are policed internally (Fiesler and Bruckman, 2019). Research into the economic impacts of LLMs on writers could also support this agenda.
Creating writer-controlled language models. Community datasets could allow writers to contribute only to projects they believed in. Such work would treat writers as experts rather than resources, and could draw on research on community archives (Flinn et al., 2009; Flinn, 2007; Huvila, 2008; Caswell et al., 2016) and archival practices for machine learning (Jo and Gebru, 2020; Desai et al., 2024). Technical work that supports smaller models that can be run locally, or models that can more easily interact with sensitive training data (Min et al., 2023), could support community models which may not have the compute power of corporate ones.
@@ Appendix A Self-Reported Demographic Data from Interviewed Writers
Table 2. Writing Genre of Interviewees
@@
The writers we spoke with also share their work via a variety of publishing or delivery methods, detailed in Table 3. As with genre, the majority of writers we spoke with (18 out of 33) also share their writing through multiple venues and means.
Table 3. Method of Delivery for Interviewees
@@
Table 4. Degree of Professionalization for Interviewees
@@
Most income comes from creative writing practice 11
Most income comes from writing-related profession 10
Most income does not come from writing 8
@@
We also noted how writers self-reported their use of LLMs; see Table 5. The vast majority (23 out of 33) were ”dabblers”—those who had tried out LLMs, but do not use them in a regular or intensive manner. A significant minority (8 out of 33) use LLMs regularly in either their paratextual or central creative writing practices.
@@ Appendix B Interview Guideline
The following guideline represents our final interview guideline. It includes core questions (solid bullet points) and potential follow-up questions (dashed bullet points). The guideline acted as a structure for the interview; not all interviews went through all follow-up questions or went through the questions in this precise order, depending on how the conversation was progressing. Note that two questions use a stimulus, which was either verbally or visually presented to interviewees to aid in the conversation.
@@ B.3. Knowledge of and attitudes towards language model training data
    • Note: Use stimulus 2